OpenStudy (keana):

Point E is located at (2, −3) and point F is located at (−2, −1). Find the y value for the point that is 3/4 the distance from point E to point F. a. 1.75 b. 0.5 c. −2 d. −1.5

ganeshie8 (ganeshie8):

heard of "section formula" before ?

OpenStudy (keana):

no

ganeshie8 (ganeshie8):

wat about "midpoint formula" ?

OpenStudy (keana):

no

ganeshie8 (ganeshie8):

just take average of "y" coordinates, that gives u midpoint

ganeshie8 (ganeshie8):

(2, −3) and (−2, −1) ^^ ^^

ganeshie8 (ganeshie8):

"y" coordinate of midpoint = (-3-1)/2 = -4/2 = -2

ganeshie8 (ganeshie8):

that lies HALF the distance between given points. but u want the point that lies at 3/4 distance between given points, right ?

OpenStudy (keana):

ok

ganeshie8 (ganeshie8):